The area of a rectangle is 30 square inches. If its length is x+2 inches, and its width is 3x-7, find the actual length.
Answer by jim_thompson5910(35256)

L*W = A

(x+2)(3x-7) = 30

3x^2-7x+6x-14 = 30

3x^2-7x+6x-14-30 = 0

3x^2-x-44 = 0

(3x+11)(x-4) = 0

3x+11=0 or x-4=0

3x=-11 or x=4

x=-11/3 or x=4

Ignore the negative solution.

So the only solution is x = 4

Length: x+2 = 4+2 = 6

So the actual length is 6 inches.